2. A long straight wire has a charge per unit length of 3.00 x 10^-12 C/m. At what distance from the wire is the electric field equal to 0.600 N/c.


1
Expert's answer
2022-03-29T12:28:11-0400

The electric field due to the linear charge

"E=\frac{\lambda}{2\pi\epsilon_0 r}"

Hence, the required distance

"r=\frac{\lambda}{2\pi\epsilon_0 E}=\frac{3.00*10^{-12}}{6.28*8.85*10^{-12}*0.600}=0.0900\:\rm m"